How much of an 18% solution of sulfuric acid should be added to 380ml of a 10% solution to obtain a 15% solution?
Please help - I'm totally stumped!

Answer by mananth(16946)

--- percent ---------------- Amount
acid type I 10 ---------------- 380 ml
acid type II 18 ---------------- x ml
Mixture 15 ---------------- 380 + x ml

10 * 380 + 18 x = 15 ( 380 + x )
3800 + 18 x = 5700 + 15 x
18 x -15 x = -3800 + 5700
3 x = 1900
/ 3
x = 633.33 ml acid type II

m.ananth@hotmail.ca